# Connect Later: Improving Fine-tuning for Robustness with Targeted Augmentations

### v1

@@ Abstract
Models trained on a labeled source domain (e.g., labeled images from wildlife camera traps) often generalize poorly when deployed on an out-of-distribution (OOD) target domain (e.g., images from new camera trap locations). In the domain adaptation setting where unlabeled target data is available, self-supervised pretraining (e.g., masked autoencoding or contrastive learning) is a promising method to mitigate this performance drop. Pretraining improves OOD error when the generic data augmentations used (e.g., masking or cropping) connect the source and target domains, which may be far apart in the input space. In this paper, we show on real-world tasks that standard fine-tuning after pretraining does not consistently improve OOD error over simply training from scratch on labeled source data. To better leverage pretraining for distribution shifts, we propose Connect Later: after pretraining with generic augmentations, fine-tune with targeted augmentations designed with knowledge of the distribution shift. Pretraining learns good representations within the source and target domains, while targeted augmentations connect the domains better during fine-tuning. Connect Later improves average OOD error over standard fine-tuning and supervised learning with targeted augmentations on 4 real-world datasets: Connect Later achieves the state-of-the-art on astronomical time-series classification (AstroClassification) by 2.5%, wildlife species identification (iWildCam-WILDS) with ResNet-50 by 0.9%, and tumor identification (Camelyon17-WILDS) with DenseNet121 by 1.1%; as well as best performance on a new dataset for astronomical time-series redshift prediction (Redshifts) by 0.03 RMSE (11% relative). [^1] Code and datasets are available at https://github.com/helenqu/connect-later.
@@ 1 Introduction
In many real-world scenarios, machine learning models are deployed on data that differ significantly from the training data [1, 2]. We focus on unsupervised domain adaptation [3, 4, 5], where we have labeled data from a source domain and unlabeled data from a target domain. We aim to learn a model that generalizes well to these out-of-distribution (OOD) target domain inputs. A real-world example is wildlife identification, where the task is to identify animal species from static camera trap images. However, human labels are only available for images from a small subset of these cameras. Images from labeled cameras may not be representative of the habitats and characteristics of unlabeled camera images.
@@
Pretraining on broad unlabeled data has shown promising results on improving OOD error in real-world problems[6, 7, 8, 9]. In particular, contrastive pretraining has been shown to learn representations that transfer well across domains [7, 10]. In contrast to conventional domain adaptation methods that focus on learning domain-invariant features [11, 12, 13, 14, 15, 16], contrastive pretraining learns representations that are not domain-invariant, but instead decompose the class and domain information, facilitating transfer across domains [7]. A favorable decomposition depends on the generic data augmentations used during contrastive pretraining to connect the source and target domains in a structured way. Intuitively, augmented (e.g. masked or cropped) source and target inputs are more likely to look similar if they are from the same class (e.g., cropping out the face of a lion in different habitats) than from different classes (e.g., no body parts of elephants and lions are alike). However, these generic augmentations must connect the domains without knowledge of the distribution shift.
@@
In this paper, we find on real-world benchmarks that standard fine-tuning after contrastive pretraining is not always effective for improving OOD error over purely supervised learning from scratch with labeled source data (Section 3). On the other hand, supervised learning with targeted augmentations [17] designed for the distribution shift improves OOD error over the supervised learning baseline on all datasets without access to any target unlabeled data. Thus, pretraining does not always learn representations that transfer across domains with standard fine-tuning.
@@
To better leverage pretraining for domain adaptation, we propose the Connect Later framework (Figure 1): after pretraining with generic augmentations, fine-tune with targeted augmentations (Section 4). Intuitively, pretraining learns good representations within each domain, while targeted augmentations better connect the domains. On a simple theoretical example where contrastive pretraining fails, Connect Later generalizes well to the target domain. We provide a general methodology for constructing these targeted augmentations, where the augmented inputs match the target distribution on a feature space where the domains differ.
@@
We evaluate our framework on 4 real-world datasets: wildlife identification [iWildCam-WILDS, 18, 9], tumor detection [Camelyon17-WILDS, 19, 9] and 2 astronomical time series tasks, AstroClassification and Redshifts, which we curate from The PLAsTiCC team et al. [20]. In Section 5, we show that Connect Later improves ID and OOD performance over standard fine-tuning or supervised learning with targeted augmentations across all datasets. Although our understanding stems from contrastive learning, we empirically apply Connect Later to better leverage pretrained representations from both masked autoencoding and contrastive learning. Connect Later achieves the state-of-the-art on three benchmarks, improving OOD accuracy on AstroClassification by 3% [21], iWildCam-WILDS with ResNet-50 by 0.9%, and Camelyon17-WILDS with DenseNet121 by 1.1%. We also contribute the Redshifts dataset, on which Connect Later produces 11% and 14% relative improvement over standard fine-tuning and supervised learning with targeted augmentations, respectively.
@@
Figure 1: Overview of the Connect Later framework for domain adaptation applied to a toy binary classification problem with two domains (filled and unfilled points), showing the representations from contrastive pretraining in $\mathbb{R}^{2}$. (Left) After contrastive pretraining with generic augmentations, the classes within each domain are linearly separable in representation space. Since the domains are far apart in input space, generic augmentations may not connect the domains, resulting in misalignment in the pretrained representations. In this case, a classifier (with a linearly extrapolating decision boundary, dashed and solid line) learned on labeled source data will misclassify the target data. (Right) Connect Later employs targeted augmentations (filled points with black border), which are designed with knowledge of the distribution shift, to connect the domains better, resulting in a classifier that generalizes well to the target domain.
@@ 2 Setup
Augmented inputs ${x^{\prime}}\in\mathcal{X}$ are drawn from an augmentation distribution $\mathcal{A}(\cdot|x)$, given an input $x\in\mathcal{X}$. Training with augmented inputs is often used to improve robustness [22, 23] and is a crucial part of the objective in contrastive pretraining [6, 7, 24]. In this work, we define two distinct augmentation distributions, $\mathcal{A}_{\text{pre}}$ and $\mathcal{A}_{\text{ft}}$, for the pretraining and fine-tuning steps, respectively. Typically, the pretraining augmentations $\mathcal{A}_{\text{pre}}$ are generic transformations, such as random cropping in vision or masking in NLP [6, 25, 26, 8, 7, 27, 24]. Fine-tuning augmentations $\mathcal{A}_{\text{ft}}$ have not been studied extensively and are typically also generic or simply the identity transformation [9, 24].
@@
Contrastive pretraining for domain adaptation consists of two steps: self-supervised pretraining on unlabeled data, then supervised fine-tuning on labeled source data [7]. For simplicity below, we consider the population objectives. Contrastive learning aims to learn an encoder which maps augmented views of the same input to similar features (“positive pairs”) and views of different inputs to dissimilar features (“negative pairs”), according to some distance metric. Formally, let $S_{+}(x,x^{+})=\mathbb{E}_{\bar{x}\sim P_{U}}[\mathcal{A}_{\text{pre}}(x\mid\bar{x})\mathcal{A}_{\text{pre}}(x^{+}\mid\bar{x})]$ be the distribution over positive pairs, which are augmentations of a single input $\bar{x}$. We pretrain an encoder $\phi:\mathcal{X}\rightarrow\mathbb{R}^{k}$ to minimize the distance $d_{+}$ between positive pair embeddings and maximize the distance $d_{-}$ between negative pair embeddings:
$\mathcal{L}_{\text{pretrain}}(\phi)=\mathbb{E}_{(x,x^{+})\sim S_{+}}[d_{+}(\phi(x),\phi(x^{+}))]-\mathbb{E}_{x,{x^{\prime}}\sim P_{U}}[d_{-}(\phi(x),\phi({x^{\prime}}))].$ (1)
@@
We refer to standard fine-tuning as the pretraining+fine-tuning procedure where $\mathcal{A}_{\text{ft}}(x^{\prime}\mid x)=1$ if $x^{\prime}=x$ (no fine-tuning augmentations). In our experiments, the standard fine-tuning procedure is linear probing then fine-tuning (LP-FT) [28], which has been shown to improve ID and OOD performance over vanilla fine-tuning. In LP-FT, we first learn a linear predictor on top of frozen pretrained features before fine-tuning all the parameters jointly.
@@ 3 Pretraining may not improve OOD performance
Table 1: Contrastive pretraining with standard fine-tuning substantially improves OOD performance for Camelyon17-WILDS but is not as effective for iWildCam-WILDS. Results are averaged over 15 trials for iWildCam-WILDS and 20 trials for Camelyon17-WILDS, and we report the 95% confidence intervals on each mean estimate.
@@
We compare ERM and standard fine-tuning on two benchmark datasets, iWildCam-WILDS (wildlife species identification) and Camelyon17-WILDS (tumor detection). In Table 1, we show that standard fine-tuning on a model pretrained using SwAV contrastive learning [6] makes minimal gains over ERM on iWildCam-WILDS ($46.4\rightarrow 46.4$ ID, $30.4\rightarrow 31.2$ OOD) compared to Camelyon17-WILDS ($89.3\rightarrow 92.3$ ID, $65.2\rightarrow 91.4$ OOD). This result runs contrary to empirical results demonstrating that contrastive pretraining is an effective domain adaptation method [6, 7, 8, 9]. We hypothesize that the generic pretraining augmentations connect the domains better for some tasks and distribution shifts than others.
@@
To understand this phenomenon, we provide a simple binary classification example of when contrastive pretraining fails for domain adaptation, following a similar augmentation graph construction to Shen et al. [7], in Appendix E. When the connectivity structure misaligns the source and target domains, such that examples from the same class are less “connected” than examples from different classes across the domains, a linear classifier trained on these pretrained representations will not transfer from source to target. This could happen, for example, when the source and target are far apart in input space and connectivity is low between examples from the same class across different domains.
@@ 3.1 Robustness gains from pretraining depend on dataset connectivity
To test this hypothesis, we empirically measure connectivity as defined in Shen et al. [7]. We follow [7] and work in the augmentation graph setting, where nodes are inputs and edge weights are the positive-pair probabilities given by $S_{+}$. We define connectivity between a class-domain pair $((y_{1},d_{1}),(y_{2},d_{2}))$ under four scenarios:
$\begin{cases}{\rho}&y_{1}=y_{2},d_{1}=d_{2}~{}\text{~{}~{}(same class, same domain)}\\ {\alpha}&y_{1}=y_{2},d_{1}\neq d_{1}\text{~{}~{}(same class, different domain)}\\ {\beta}&y_{1}\neq y_{2},d_{1}=d_{2}\text{~{}~{}(different class, same domain)}\\ {\gamma}&y_{1}\neq y_{2},d_{1}\neq d_{2}\text{~{}~{}(different class and domain)}\\ \end{cases},$ (3)
where each value is an average edge weight over the edges that satisfy each case. Shen et al. [7] show in simple augmentation graphs that contrastive pretraining theoretically learns transferable representations when ${\alpha}>{\gamma}$ and ${\beta}>{\gamma}$, and that the ratios $\frac{{\alpha}}{{\gamma}}$ and $\frac{{\beta}}{{\gamma}}$ empirically correlate well with OOD accuracy. Intuitively, the pretraining augmentations are less likely to change both the domain and class of an input than changing just domain or just class.
Table 2: Empirically estimated connectivity measures for iWildCam-WILDS and Camelyon17-WILDS. From Shen et al. [7], contrastive pretraining theoretically learns transferable representations for UDA when both across-domain ($\alpha$) and across-class ($\beta$) connectivity is greater than across-both ($\gamma$), using notation from  [7]. In iWildCam-WILDS, across-both connectivity $>$ across-class, which violates the condition, while Camelyon17-WILDS satisfies the condition.
across-domain ($\alpha$) across-class ($\beta$) across-both ($\gamma$)
@@
To better understand why contrastive pretraining performs differently on iWildCam-WILDS and Camelyon17-WILDS, we empirically evaluate the connectivity measures for iWildCam-WILDS and Camelyon17-WILDS, following Shen et al. [7]. Using augmented inputs from 2 class-domain pairs, we train a binary classifier to predict the class-domain pair of each input, and interpret the test error of the classifier as an estimate for connectivity. We average each connectivity value over 15 class-domain pairs (see Appendix D for details). Our results, summarized in Table 2, show that iWildCam-WILDS connectivity measures violate the condition for contrastive pretraining in the UDA setting, since across-both connectivity $>$ across-class ($\gamma>\beta$). This finding is consistent with our observation that contrastive pretraining fails for iWildCam-WILDS while producing gains for Camelyon17-WILDS, and further underscores the need for domain adaptation methods that correct the misaligned connectivity structure.

### v2

@@ Abstract
Models trained on a labeled source domain often generalize poorly when deployed on an out-of-distribution (OOD) target domain. In the domain adaptation setting where unlabeled target data is available, self-supervised pretraining (e.g., contrastive learning or masked autoencoding) is a promising method to mitigate this performance drop. Pretraining depends on generic data augmentations (e.g., cropping or masking) to learn representations that generalize across domains, which may not work for all distribution shifts. In this paper, we show on real-world tasks that standard fine-tuning after pretraining does not consistently improve OOD error over simply training from scratch on labeled source data. To better leverage pretraining for distribution shifts, we propose the Connect Later framework, which fine-tunes the model with targeted augmentations designed with knowledge of the shift. Intuitively, pretraining learns good representations within the source and target domains, while fine-tuning with targeted augmentations improves generalization across domains. Connect Later achieves state-of-the-art OOD accuracy while maintaining comparable or better in-distribution accuracy on 4 real-world tasks in wildlife identification (iWildCam-WILDS), tumor detection (Camelyon17-WILDS), and astronomy (AstroClassification, Redshifts).
robustness, machine learning, pretraining, fine-tuning
@@ 1 Introduction
In the real world, machine learning models are often deployed on data that differ significantly from training data (Quiñonero-Candela et al., 2009; Koh et al., 2021). We focus on unsupervised domain adaptation (Shimodaira, 2000; Blitzer et al., 2006; Sugiyama et al., 2007), where we have labeled data from a source domain and unlabeled data from a target domain. We aim to learn a model that generalizes well to these out-of-distribution (OOD) target domain inputs. A real-world example is wildlife identification, where the task is to identify animal species from static camera trap images. However, human labels are only available for images from a small subset of these cameras, which may not be representative of the habitats and characteristics of unlabeled camera images.
@@
Pretraining on broad unlabeled data has shown promising results on improving OOD error in real-world problems (Caron et al., 2020; Shen et al., 2022; Radford et al., 2021; Sagawa et al., 2022). In particular, contrastive pretraining has been shown to learn representations that transfer well across domains (Shen et al., 2022; HaoChen et al., 2022). In contrast to conventional domain adaptation methods that focus on learning domain-invariant features (Ganin et al., 2016; Kang et al., 2019; Tzeng et al., 2017; Saenko et al., 2010; Sun et al., 2016; Hoffman et al., 2018), contrastive pretraining learns representations that are not domain-invariant, but instead decompose the class and domain information, facilitating transfer across domains (Shen et al., 2022). A favorable decomposition depends on the generic data augmentations used during contrastive pretraining to align representations across domains. Intuitively, augmented (e.g. masked or cropped) source and target inputs should be more likely to look similar if they are from the same class (e.g., cropping out the face of a lion in different habitats) than from different classes (e.g., no body parts of elephants and lions are alike). However, these generic augmentations may not be suitable for all distribution shifts.
@@
In this paper, we find on real-world benchmarks that standard fine-tuning after contrastive pretraining is not always effective for improving OOD error over purely supervised learning from scratch with labeled source data (Section 3). On the other hand, supervised learning with targeted augmentations (Gao et al., 2023) designed for the distribution shift improves OOD error over the supervised learning baseline on all datasets without access to any target unlabeled data. Thus, pretraining does not always learn representations that transfer across domains with standard fine-tuning.
@@
To better leverage pretraining for domain adaptation, we propose the Connect Later framework (Figure 1): after pretraining with generic augmentations, fine-tune with targeted augmentations (Section 4). Intuitively, pretraining learns good representations within each domain, while targeted augmentations incorporate domain knowledge to improve generalization across domains. Through both empirical and theoretical examples, we show that Connect Later generalizes well to the target domain even in scenarios where pretraining alone produces minimal OOD performance improvements. We provide a general methodology for constructing these targeted augmentations by matching augmented inputs to the target distribution on a feature space where the domains differ.
@@
We evaluate our framework on 4 real-world datasets: wildlife identification (iWildCam-WILDS, Beery et al., 2020; Sagawa et al., 2022), tumor detection (Camelyon17-WILDS, Bandi et al., 2018; Sagawa et al., 2022) and 2 astronomical time series tasks, AstroClassification and Redshifts, which we curate from The PLAsTiCC team et al. (2018). In Section 5, we show that Connect Later improves OOD performance over standard fine-tuning or supervised learning with targeted augmentations across all datasets. Although our understanding stems from contrastive learning, we empirically apply Connect Later to better leverage pretrained representations from both masked autoencoding and contrastive learning. Connect Later achieves the state-of-the-art on three benchmarks, improving OOD accuracy on AstroClassification by 3% (Boone, 2019), iWildCam-WILDS with ResNet-50 by 0.9%, and Camelyon17-WILDS with DenseNet121 by 1.1%. We also contribute the Redshifts dataset, on which Connect Later outperforms the best baseline by 11% relative improvement.
@@
Figure 1: Overview of the Connect Later framework applied to a toy binary classification problem with two domains (filled and unfilled points), showing the representations from contrastive pretraining in $\mathbb{R}^{2}$. (Left) After contrastive pretraining with generic augmentations, the classes within each domain are linearly separable in representation space. Since the domains are far apart in input space, generic augmentations may misalign the pretrained representations. In this case, a classifier (with a linearly extrapolating decision boundary, dashed and solid line) learned on labeled source data will misclassify the target data. (Right) Connect Later employs targeted augmentations (filled points with black border) designed with knowledge of the distribution shift to connect the domains better, resulting in a classifier that generalizes well to the target domain.
@@ 2 Setup
Augmented inputs ${x^{\prime}}\in\mathcal{X}$ are drawn from an augmentation distribution $\mathcal{A}(\cdot|x)$, given an input $x\in\mathcal{X}$. Training with augmented inputs is often used to improve robustness (Hendrycks et al., 2019, 2020) and is crucial to contrastive pretraining (Caron et al., 2020; Shen et al., 2022; Devlin et al., 2019). In this work, we define two distinct augmentation distributions, $\mathcal{A}_{\text{pre}}$ and $\mathcal{A}_{\text{ft}}$, for the pretraining and fine-tuning steps, respectively. Typically, the pretraining augmentations $\mathcal{A}_{\text{pre}}$ are generic transformations, such as random cropping in vision or masking in NLP (Caron et al., 2020; Chen et al., 2020; He et al., 2020; Radford et al., 2021; Shen et al., 2022; He et al., 2022; Devlin et al., 2019). Fine-tuning augmentations $\mathcal{A}_{\text{ft}}$ have not been studied extensively and are typically also generic or simply the identity transformation (Sagawa et al., 2022; Devlin et al., 2019).
@@
Contrastive pretraining for domain adaptation consists of two steps: self-supervised pretraining on unlabeled data, then supervised fine-tuning on labeled source data (Shen et al., 2022). For simplicity below, we consider the population objectives. Contrastive learning aims to learn an encoder which maps augmented views of the same input to similar features (“positive pairs”) and views of different inputs to dissimilar features (“negative pairs”), according to some distance metric. Formally, let $S_{+}(x,x^{+})=\mathbb{E}_{\bar{x}\sim P_{U}}[\mathcal{A}_{\text{pre}}(x\mid\bar{x})\mathcal{A}_{\text{pre}}(x^{+}\mid\bar{x})]$ be the distribution over positive pairs, which are augmentations of a single input $\bar{x}$. We pretrain an encoder $\phi:\mathcal{X}\rightarrow\mathbb{R}^{k}$ to minimize the distance $d_{+}$ between positive pair embeddings and maximize the distance $d_{-}$ between negative pair embeddings:
$\begin{split}\mathcal{L}_{\text{pretrain}}(\phi)=\mathbb{E}_{(x,x^{+})\sim S_{+}}[d_{+}(\phi(x),\phi(x^{+}))]-\\ \mathbb{E}_{x,{x^{\prime}}\sim P_{U}}[d_{-}(\phi(x),\phi({x^{\prime}}))].\end{split}$ (1)
@@
We refer to standard fine-tuning as the pretraining+fine-tuning procedure where $\mathcal{A}_{\text{ft}}(x^{\prime}\mid x)=1$ if $x^{\prime}=x$ (no fine-tuning augmentations). In our experiments, the standard fine-tuning procedure is linear probing then fine-tuning (LP-FT) (Kumar et al., 2022), which has been shown to improve ID and OOD performance over vanilla fine-tuning. In LP-FT, we first learn a linear predictor on top of frozen pretrained features before fine-tuning all the parameters jointly.
@@ 3 Pretraining may not improve OOD performance
Table 1: Contrastive pretraining with standard fine-tuning substantially improves OOD performance for Camelyon17-WILDS but is not very effective for iWildCam-WILDS. Results are averaged over 15 trials for iWildCam-WILDS and 20 trials for Camelyon17-WILDS, and we report the 95% confidence intervals on each mean estimate.
@@
We compare ERM and standard fine-tuning on two benchmark datasets, iWildCam-WILDS (wildlife species identification) and Camelyon17-WILDS (tumor detection). In Table 1, we show that standard fine-tuning on a model pretrained using SwAV contrastive learning (Caron et al., 2020) makes minimal gains over ERM on iWildCam-WILDS ($46.4\rightarrow 46.4$ ID, $30.4\rightarrow 31.2$ OOD) compared to Camelyon17-WILDS ($89.3\rightarrow 92.3$ ID, $65.2\rightarrow 91.4$ OOD). This result runs contrary to prior work demonstrating that contrastive pretraining is an effective domain adaptation method (Caron et al., 2020; Shen et al., 2022; Radford et al., 2021; Sagawa et al., 2022). We hypothesize that the generic pretraining augmentations connect the domains better for some tasks and distribution shifts than others.
@@
To understand this phenomenon, we provide a simple binary classification example of when contrastive pretraining fails for domain adaptation, following a similar augmentation graph construction to Shen et al. (2022), in Appendix E. When the connectivity structure misaligns the source and target domains, such that examples from the same class are less “connected” than examples from different classes across the domains, a linear classifier trained on these pretrained representations will not transfer from source to target. This could happen, for example, when the source and target are far apart in input space and connectivity is low between examples from the same class across different domains.
@@ 3.1 Robustness gains from pretraining depend on dataset connectivity
To better understand why contrastive pretraining performs differently on these two datasets, we empirically evaluate the connectivity measures for iWildCam-WILDS and Camelyon17-WILDS. We follow Shen et al. (2022) and work in the augmentation graph setting, where nodes are inputs and edge weights are the positive-pair probabilities given by $S_{+}$. We define connectivity between a class-domain pair $((y_{1},d_{1}),(y_{2},d_{2}))$ under four scenarios:
$\begin{cases}{\rho}&y_{1}=y_{2},d_{1}=d_{2}~{}\text{~{}~{}(same class, same domain)}\\ {\alpha}&y_{1}=y_{2},d_{1}\neq d_{2}\text{~{}~{}(same class, different domain)}\\ {\beta}&y_{1}\neq y_{2},d_{1}=d_{2}\text{~{}~{}(different class, same domain)}\\ {\gamma}&y_{1}\neq y_{2},d_{1}\neq d_{2}\text{~{}~{}(different class and domain)}\\ \end{cases},$ (3)
where each value is an average edge weight over the edges that satisfy each case. Shen et al. (2022) show in simple augmentation graphs that contrastive pretraining theoretically learns transferable representations when ${\alpha}>{\gamma}$ and ${\beta}>{\gamma}$, and that the ratios $\frac{{\alpha}}{{\gamma}}$ and $\frac{{\beta}}{{\gamma}}$ empirically correlate well with OOD accuracy. Intuitively, the pretraining augmentations are less likely to change both the domain and class of an input than changing just domain or just class.
Table 2: Empirically estimated connectivity measures for iWildCam-WILDS and Camelyon17-WILDS. From Shen et al. (2022), contrastive pretraining theoretically learns transferable representations for UDA when both across-domain ($\alpha$) and across-class ($\beta$) connectivity is greater than across-both ($\gamma$). In iWildCam-WILDS, $\gamma>\beta$, violating the condition, while Camelyon17-WILDS satisfies the condition.
$\alpha$ $\beta$ $\gamma$
@@
We empirically evaluate the connectivity measures for iWildCam-WILDS and Camelyon17-WILDS following Shen et al. (2022). Using augmented inputs from 2 class-domain pairs, we train a binary classifier to predict the class-domain pair of each input, and interpret the test error of the classifier as an estimate for connectivity. We average each connectivity value over 15 class-domain pairs (see Appendix D for details). Our results, summarized in Table 2, show that iWildCam-WILDS connectivity measures violate the condition for contrastive pretraining in the UDA setting, since across-both connectivity $>$ across-class ($\gamma>\beta$). This finding is consistent with our observation that contrastive pretraining is far less effective for iWildCam-WILDS compared to Camelyon17-WILDS, and further underscores the need for domain adaptation methods that correct the misaligned connectivity structure.